# Large Language Model Sentinel: LLM Agent for Adversarial Purification

### v3

@@ Abstract
Over the past two years, the use of large language models (LLMs) has advanced rapidly. While these LLMs offer considerable convenience, they also raise security concerns, as LLMs are vulnerable to adversarial attacks by some well-designed textual perturbations. In this paper, we introduce a novel defense technique named Large LAnguage MOdel Sentinel (LLAMOS), which is designed to enhance the adversarial robustness of LLMs by purifying the adversarial textual examples before feeding them into the target LLM. Our method comprises two main components: a) Agent instruction, which can simulate a new agent for adversarial defense, altering minimal characters to maintain the original meaning of the sentence while defending against attacks; b) Defense guidance, which provides strategies for modifying clean or adversarial examples to ensure effective defense and accurate outputs from the target LLMs. Remarkably, the defense agent demonstrates robust defensive capabilities even without learning from adversarial examples. Additionally, we conduct an intriguing adversarial experiment where we develop two agents, one for defense and one for attack, and engage them in mutual confrontation. During the adversarial interactions, neither agent completely beat the other. Extensive experiments on both open-source and closed-source LLMs demonstrate that our method effectively defends against adversarial attacks, thereby enhancing adversarial robustness.
Introduction
Figure 1: Illustration of adversarial attacks and adversarial purification on large language models. (a) When the clean example $x$ is input into the target LLM, a correct label will be output. However, after the attack, the adversarial example $x^{\prime}$ will make the target LLM outputs an incorrect label. We use LLAMOS (b) to purify the adversarial example before inputting it into the target LLM, ensuring that the purified example $\hat{x}$ can be predicted as the correct label.
@@ Introduction
Large language models (LLMs) have garnered significant attention due to their impressive performance across a wide range of natural language tasks (Minaee et al. 2024). The pre-trained LLMs, such as Meta’s LLAMA (Touvron et al. 2023a, b) and OpenAI’s ChatGPT (OpenAI 2022; Achiam et al. 2023), have become essential foundations for AI applications in various sectors such as healthcare, education, and visual tasks (Kasneci et al. 2023; Thirunavukarasu et al. 2023; OpenAI 2023; Köpf et al. 2024; Romera-Paredes et al. 2024). Despite their widespread use and convenience, concerns about the security of these models are increasing. Specifically, LLMs have been shown to be vulnerable to adversarial textual examples (Wang et al. 2023a; Xu et al. 2024), which involve subtle modifications to textual content that maintain the same meaning for humans but completely change the prediction results to LLMs, often with severe consequences.
@@
To achieve robust defense against adversarial attacks on LLMs, a prevalent strategy is fine-tuning the LLMs with adversarial examples to enhance model alignment (Shen et al. 2023; Wang et al. 2023c). LLM-based adversarial fine-tuning (AFT) can be implemented either through in-context learning (Dong et al. 2022; Xiang et al. 2024) or by optimizing the parameters of pre-trained LLMs using adversarial examples (Dettmers et al. 2024; Li et al. 2024b). However, LLM-based AFT methods necessitate additional computational resources and training time. Achieving robust and reliable LLMs typically requires significant costs (Hu et al. 2022), which is prohibitive for ordinary users. Additionally, due to the discrete nature of textual information, these adversarial examples can have substitutes in any token of the sentence, with each having a large candidate list (Li, Song, and Qiu 2023). This leads to a combinatorial explosion, making the application of AFT methods challenging or resulting in poor generalization when trained on a limited dataset of adversarial examples. Consequently, developing an efficient and user-friendly robust LLM system remains a huge challenge and an urgent issue that continues to be addressed.
@@
We comprehensively evaluate the performance of our method on GLUE datasets, conducting experiments with both representative open-source and closed-source LLMs, LLAMA-2 (Touvron et al. 2023b) and GPT-3.5 (OpenAI 2022). The experimental results demonstrate that the LLM-based AP method effectively defends against adversarial attacks. Specifically, our proposed method achieves a maximum reduction in the attack success rate (ASR) by up to 45.59% and 37.86% with LLAMA-2 and GPT-3.5, respectively. Additionally, we observe that the initial defense agent fails to achieve the expected results under some obvious attacks. Therefore, we employ the in-context learning (Dong et al. 2022) to further optimize the defense agent, significantly enhancing the defense capabilities almost without adding any additional costs. Finally, we conduct an intriguing online adversarial experiment, creating an adversarial system using two LLM-based agents (one for defense and one for attack) along with a target LLM for classification. During the adversarial interaction, the defense agent and attack agent continuously counter each other, resembling the adversarial training process in traditional image tasks (Goodfellow, Shlens, and Szegedy 2015).
Our contributions are summarized as follows.
  • We propose a novel defense technique named LLAMOS, which aims to purify the adversarial textual examples before feeding them into the target LLM. To the best of our knowledge, we are the first to employ an LLM agent to enhance the adversarial robustness of LLMs.
  • The defense agent in LLAMOS is a plug-and-play module, serving as a pre-processing step. Notably, it operates without retraining of target LLM, rendering it efficient and user-friendly.
  • We conduct extensive experiments to empirically demonstrate that the proposed method can effectively defend against adversarial attacks.
Preliminary
This section briefly reviews the adversarial attacks and evaluations of adversarial robustness on LLMs.
Adversarial Attcks on LLMs
@@
Given a target LLM $f_{t}$ with task instruction, input $x$, and correct output $y$, the adversarial attacks aim to find the adversarial examples $x^{\prime}$ that can fool the target LLM $f_{t}$ on classification tasks. The adversarial examples $x^{\prime}$ can be obtained by the LLM itself $f_{atk}$ with different system prompt (Xu et al. 2024),
@@
Evaluations of LLMs Robustness
To evaluate the effectiveness of the defense method, we follow the setting from Wang et al. (2021); Xu et al. (2024), using the attack success rate (ASR) and traditional robust accuracy (RA) on the adversarial examples as measures of the robustness of the defense method.
$\text{ASR}=\frac{\sum_{(x,y)\in D}\mathds{1}\left[f_{t}(x^{\prime})\neq y\right]\cdot\mathds{1}\left[f_{t}(x)=y\right]}{\sum_{(x,y)\in D}\mathds{1}\left[f_{t}(x)=y\right]},$ (1)
$\text{RA}=\frac{1}{N}\sum_{(x^{\prime},y)\in D^{\prime}}\mathds{1}\left[f_{t}(x^{\prime})=y\right],$ (2)
where $\mathds{1}[\cdot]\in\{0,1\}$ is an indicator function. $D$ is the original test dataset and $D^{\prime}$ is the adversarial example dataset, and $N$ is the number of examples. The lower the ASR, the higher the RA, indicating greater model robustness.
Methods
We propose a novel defense technique for large language model-based adversarial purification (LLAMOS), which purifies adversarial examples by the LLM-based defense agent before feeding examples into the target LLM. Firstly, we outline the overall pipeline of LLAMOS. Subsequently, we further augment the defense agent using in-context learning. Finally, we present the design of the adversarial system, incorporating the defense agent, attack agent, and target LLM.
Overview of LLAMOS
@@ Methods
In this paper, we utilize the existing LLMs denoted by target LLMs $f_{t}$ as classifiers. Given an adversarial example $x^{\prime}$, the target LLM $f_{t}$ outputs the incorrect label $y^{\prime}=f_{t}(x^{\prime})$, while after the purification, the purified example $\hat{x}=g_{stl}(x^{\prime})$ can be predicted as the correct label $y=f_{t}(g_{stl}(x^{\prime}))$. To achieve this, we design prompts for generating a defense agent $g_{stl}$ as described in the following.
@@
You can complete the task using the following guidance: [Defense Guidance].
Input: [Input]. Now, let’s start the defense process and only output the generated sentence.
Input Description. The format of Input $x_{in}\in\{D_{i},i=1...6\}$ varies significantly across different datasets $D_{i}$, necessitating tailored input formats $x_{in}$ to correspond with the specific structure and content of each dataset, details in Table 10. For instance, the SST-2 dataset (Socher et al. 2013) typically consists of a single sentence per data point. On the other hand, the MNLI dataset (Williams, Nangia, and Bowman 2018) is structured to include pairs of sentences labeled as premise and hypothesis.
Task Description. Similar to the input descriptions, the tasks associated with each dataset $D_{i}$ are distinct. As illustrated earlier, SST-2 focuses on determining the sentiment of a given sentence, making it a straightforward classification challenge. Conversely, MNLI presents a more complex task of natural language inference, where the relationship between a pair of sentences must be discerned and classified correctly. Detailed input and task descriptions are provided in Table 9111Tables 9 to 12 are in the Appendix..
Defense Goal. Based on traditional adversarial purification methods (Shi, Holtz, and Mishne 2021; Srinivasan et al. 2021; Lin et al. 2024b, c), we design the goal for LLAMOS as follows: “Keeping the semantic meaning of the new sentence the same as the original one; For natural examples, the new sentence should remain unchanged. For adversarial examples, modify the sentence so that it is classified as the correct label, effectively reversing the adversarial effect.”
Defense Guidance. The defense guidance offers specific instructions to the defense agent on how to modify the input text to ensure effective defense and accurate outputs from the target LLM. In designing our guidance, we considered attacks at various levels (Xu et al. 2024), including character, word, and sentence levels, which are presented in Table 1. These guidances are not rigidly fixed; they can be fine-tuned according to specific tasks.
Table 1: The defense guidances for defense agent.
      Index [Defense Guidance]
@@
      4 If necessary, feel free to replace, delete, add words, or adjust the word order.
@@
After the defense agent generates the new sentence, the purified example $\hat{x}=g_{stl}(x^{\prime})$ is input into the target LLM $f_{t}$ for classification.
Enhencing the Defense with In-Context Learning
In the initial defense agent, the defense guidance relies on common sense, which may result in poor performance against some special attacks, even when the attacker adds obvious characters. To address this limitation, we introduce ICL (Dong et al. 2022) to further optimize the defense agent by,
@@
The new sentence still contains a lot of harmful content caused by adversarial attacks, such as [Specific Guidance]. Please consider these contents and output a new sentence for me.
Input: [Input]. Now, let’s start the defense process and only output the generated sentence.
The specific guidance is designed to assist the defense agent in better understanding an attack and generating a new sentence capable of effectively defending against the attack. These guidelines can be fine-tuned to address specific attacks and can be incorporated into the defense agent as needed. Through in-context learning, the defense agent can be continuously optimized, significantly enhancing its performance almost without adding any additional costs.
Adversarial System with Multiple LLMs
In this section, we devise an adversarial system involving multiple LLMs. Given that our method introduces a defense agent against attackers, a natural idea is to then create an attack agent to counter the defender. The attack agent is tasked with generating adversarial examples from purified examples to deceive the target LLM once more. To accomplish this, we design prompts for generating an attack agent $g_{atk}$ by,
{mdframed}
# Attack Agent Instruction
@@
For example, the original sentence [Purified Example] is classified as [Correct Label]. You should generate a new sentence which is classified as [Incorrect Label].
Input: [Input]. Now, let’s start the attack process and only output the generated sentence.
The prompt structure of the attack agent and the defense agent is basically the same, although there are some differences in details. The input description of the attack agent includes the correct label $y$, and the input format is $(x_{in},y)\in D$. The attack instruction is “1. The new sentence should be classified as the opposite of the ‘correct label’. 2. Change at most two letters in the sentence.” Finally, we provide a specific example to help the attack agent better understand the attack task.
Then, we combine the defense agent and attack agent to form an adversarial system, as illustrated in Figure 2. In the adversarial system, the purified examples can be attacked again by the attack agent, and likewise, the adversarial examples can also be purified by the defense agent. They continuously counter each other, much like adversarial training (Goodfellow, Shlens, and Szegedy 2015).
Figure 2: Adversarial system with multiple LLMs, including the defense agent, attack agent, and target LLM.
Table 2: The attack success rate (ASR) defense against PromptAttack-EN and PromptAttack-FS-EN on the GLUE dataset with GPT-3.5. The lower the ASR, the greater the model robustness.

### v4

@@ Abstract
Over the past two years, the use of large language models (LLMs) has advanced rapidly. While these LLMs offer considerable convenience, they also raise security concerns, as LLMs are vulnerable to adversarial attacks by the well-designed textual perturbations. In this paper, we introduce a novel defense technique named Large LAnguage MOdel Sentinel (LLAMOS), which is designed to enhance the adversarial robustness of LLMs by purifying the adversarial textual examples before feeding them into the target LLM. Our method comprises two main components: a) Agent instruction, which can simulate a new agent for adversarial defense, altering minimal characters to maintain the original meaning of the sentence while defending against attacks; b) Defense guidance, which provides strategies for modifying clean or adversarial examples to ensure effective defense and accurate outputs from the target LLMs. Remarkably, the defense agent demonstrates robust defensive capabilities even without learning from adversarial examples. Additionally, we conduct an intriguing adversarial experiment where we develop two agents, one for defense and one for attack, and engage them in mutual confrontation. During the adversarial interactions, neither agent completely beat the other. Extensive experiments on both open-source and closed-source LLMs demonstrate that our method effectively defends against adversarial attacks, thereby enhancing adversarial robustness.
Keywords:
Adversarial Robustness Adversarial Purification Large Language Model (LLM) LLM Agent.
1 Introduction
Figure 1: Illustration of adversarial attack and purification on LLMs. (a) The clean example $x$ is perturbed to the adversarial example $x^{\prime}$, causing misclassification. (b) LLAMOS purifies $x^{\prime}$ to the purified example $\hat{x}$, restoring the correct label.
@@ 1 Introduction
Large language models (LLMs) have garnered significant attention due to their impressive performance across a wide range of natural language tasks [18]. The pre-trained LLMs, such as Meta’s LLAMA [31, 32] and OpenAI’s ChatGPT [21], have become essential foundations for AI applications in various sectors such as healthcare, education, and visual tasks [22, 7, 24]. Despite their widespread use and convenience, concerns about the security of these models are increasing. Specifically, LLMs have been shown to be vulnerable to adversarial textual examples [34, 41], which involve subtle modifications to textual content that maintain the same meaning for humans but completely change the prediction results to LLMs, often with severe consequences.
@@
To achieve robust defense against adversarial attacks on LLMs, a prevalent strategy is fine-tuning the LLMs with adversarial examples to enhance model alignment [25, 37]. LLM-based adversarial fine-tuning (AFT) can be implemented either through in-context learning [4, 40] or by optimizing the parameters of pre-trained LLMs using adversarial examples [3, 11]. However, LLM-based AFT methods necessitate additional computational resources and training time. Achieving robust and reliable LLMs typically requires significant costs [6], which is prohibitive for ordinary users. Additionally, due to the discrete nature of textual information, these adversarial examples can have substitutes in any token of the sentence, with each having a large candidate list [9]. This leads to a combinatorial explosion, making the application of AFT methods challenging or resulting in poor generalization when trained on a limited dataset of adversarial examples. Consequently, developing an efficient and user-friendly robust LLM system remains a huge challenge and an urgent issue that continues to be addressed.
@@
We comprehensively evaluate the performance of our method on GLUE datasets, conducting experiments with both representative open-source and closed-source LLMs, LLAMA-2 [32] and GPT-3.5 [21]. The experimental results demonstrate that the LLM-based AP method effectively defends against adversarial attacks. Specifically, our proposed method achieves a maximum reduction in the attack success rate (ASR) by up to 45.59% and 37.86% with LLAMA-2 and GPT-3.5, respectively. Additionally, we observe that the initial defense agent fails to achieve the expected results under some obvious attacks. Therefore, we employ the in-context learning [4] to further optimize the defense agent, significantly enhancing the defense capabilities almost without adding any additional costs. Finally, we conduct an intriguing online adversarial experiment, creating an adversarial system using two LLM-based agents (one for defense and one for attack) along with a target LLM for classification.
2 Related Work
2.0.1 Adversarial Attack
Deep neural networks (DNNs) are vulnerable to adversarial examples [30], which are generated by adding small, human-imperceptible perturbations to natural examples, but completely change the prediction results to DNNs [5, 13]. With the rapidly increasing applications of LLMs [22, 7, 24], security concerns have emerged as a critical area of research [9, 43], with researchers increasingly focusing on adversarial attacks targeting LLMs. In a similar setup to DNNs, for LLMs, attackers manipulate a small amount of text to change the output of the target LLM while maintaining the semantic information for humans [35, 41]. Presently, addressing the security issues surrounding LLMs is of paramount importance and requires urgent attention.
2.0.2 Adversarial Defense
There are two main defense techniques on traditional DNNs, including adversarial training (AT) [5] and adversarial purification (AP) [26]. Unlike traditional DNNs, retraining LLMs is nearly impossible due to cost issues [9]. Therefore, most methods enhance the robustness of LLMs through adversarial fine-tuning (AFT) [40, 11]. While AFT can effectively defend against attacks, it remains susceptible to unseen attacks whose adversarial examples that the LLMs have not previously learned [9]. Additionally, even with fine-tuning, training the LLMs will still consume a significant cost [6, 3]. Adversarial purification (AP) aims to purify adversarial examples before feeding them into the target model, which has emerged as a promising defense method [26, 13]. Compared with the AT or AFT method, the AP method utilizes an additional model that can defend against unseen attacks without retraining the target model [13, 9]. In some traditional computer vision and natural language processing tasks, researchers have started using LLMs for adversarial purification [27, 10, 19], but the security issues of LLMs themselves have not been deeply considered. Therefore, we propose a novel LLM defense technique named LLAMOS to purify the adversarial textual examples before feeding them into the target LLM, aiming to improve the robustness.
2.0.3 Large Language Model Agent
The LLM agent is a new research direction that has emerged in recent years [20, 17]. This novel type of agent is capable of interacting with humans in natural language, leading to a significant increase in applications across fields such as chatbots, natural sciences, robotics, and workflows [12, 15]. Furthermore, LLMs have demonstrated promising zero-shot/few-shot planning and reasoning capabilities across various configurations [29], covering specific environments and reasoning tasks [42]. In this paper, we introduce a new variant of the LLM agent designed specifically to purify adversarial textual examples generated by attacks.
3 Preliminary
3.0.1 Adversarial Attcks on LLMs
@@
Given a target LLM $f_{t}$ with task instruction, input $x$, and correct label $y$, attackers aim to find the adversarial examples $x^{\prime}$ that can fool the target LLM $f_{t}$ on classification tasks. The adversarial examples $x^{\prime}$ can be obtained by the LLM itself $f_{atk}$ with different system prompt [41],
@@
3.0.2 Evaluations of LLMs Robustness
To evaluate the effectiveness of the defense method, we follow the setting from [36, 41], using the attack success rate (ASR) and traditional robust accuracy (RA) on the adversarial examples as measures of the robustness of the defense method.
$\text{ASR}=\frac{\sum_{(x,y)\in D}\mathbbm{1}\left[f_{t}(x^{\prime})\neq y\right]\cdot\mathbbm{1}\left[f_{t}(x)=y\right]}{\sum_{(x,y)\in D}\mathbbm{1}\left[f_{t}(x)=y\right]},\quad\text{RA}=\frac{1}{N}\sum\mathbbm{1}\left[f_{t}(x^{\prime})=y\right],$ (1)
where $\mathbbm{1}[\cdot]\in\{0,1\}$ is an indicator function. $D$ is the original test dataset and $D^{\prime}$ is the adversarial example dataset, and $N$ is the number of examples. The lower the ASR, the higher the RA, indicating greater model robustness.
4 Methods
We propose a novel defense technique for large language model-based adversarial purification (LLAMOS), which purifies adversarial textual examples by the LLM-based defense agent before feeding them into the target LLM. Firstly, we outline the overall pipeline of LLAMOS. Subsequently, we further augment the defense agent using in-context learning. Finally, we present the design of the adversarial system, incorporating the defense agent, attack agent, and target LLM.
4.1 Overview of LLAMOS
@@
In this paper, we utilize the existing LLM denoted by target LLM $f_{t}$ as the classifier. Given an adversarial example $x^{\prime}$, the target LLM $f_{t}$ outputs the incorrect label $y^{\prime}=f_{t}(x^{\prime})$, while after purification, the purified example $\hat{x}=g_{llm}(x^{\prime})$ can be predicted as the correct label $y=f_{t}(g_{llm}(x^{\prime}))$. To achieve this, we design prompts for generating a defense agent $g_{llm}$ as described in the following.
@@
You can complete the task using the following guidance: [Defense Guidance]. Input: [Input]. Now, let’s start the defense process and only output the generated sentence.
Input Description: The format of Input $x_{in}\in\{D_{i},i=1...6\}$ varies significantly across different datasets $D_{i}$, necessitating tailored input formats $x_{in}$ to correspond with the specific structure and content of each dataset, details in Table 10. For instance, the SST-2 dataset [28] typically consists of a single sentence per data point. On the other hand, the MNLI dataset [39] is structured to include pairs of sentences labeled as premise and hypothesis.
Task Description: Similar to the input description, the tasks associated with each dataset $D_{i}$ are distinct. As illustrated earlier, SST-2 focuses on determining the sentiment of a given sentence, making it a straightforward classification task. Conversely, MNLI presents a more complex task of natural language inference, where the relationship between a pair of sentences must be discerned and classified correctly. Detailed input and task descriptions are provided in Table 9.
Defense Goal: Based on traditional adversarial purification [26, 13, 14], we design the goal for LLAMOS as follows: “Keeping the semantic meaning of the new sentence the same as the original one; For natural examples, the new sentence should remain unchanged. For adversarial examples, modify the sentence so that it is classified as the correct label, effectively reversing the adversarial effect.”
Defense Guidance: The defense guidance offers specific instructions to the defense agent on how to modify the input text to ensure effective defense and accurate outputs from the target LLM. In designing our guidance, we considered attacks at various levels [41], including character, word, and sentence levels, which are presented in Table 1. These guidances are not rigidly fixed; they can be fine-tuned according to specific tasks.
After the defense agent generates the new sentence, the purified example $\hat{x}=g_{llm}(x^{\prime})$ is input into the target LLM $f_{t}$ for classification.
Figure 2: Illustration of adversarial system with attack and defense agents.
Table 1: The guidances for defense agent.
Idx. [Defense Guidance]
@@
4 If necessary, feel free to replace, delete,
add words, or adjust the word order.
@@ 4 If necessary, feel free to replace, delete,
4.2 Enhencing the Defense with In-Context Learning
In the initial defense agent, the defense guidance relies on common sense, which may result in poor performance against some special attacks, even when the attacker adds obvious characters. To address this limitation, we introduce in-context learning (ICL) [4] to further optimize the defense agent by,
@@ 4.2 Enhencing the Defense with In-Context Learning
The new sentence still contains a lot of harmful content caused by adversarial attacks, such as [Specific Guidance]. Please consider these contents and output a new sentence for me. Input: [Input]. Now, let’s start the defense process and only output the generated sentence.
The specific guidance is designed to assist the defense agent in better understanding an attack and generating a new sentence capable of effectively defending against the attack. These guidances can be fine-tuned to address specific attacks and can be incorporated into the defense agent as needed. Through in-context learning, the defense agent can be continuously optimized, significantly enhancing its performance almost without adding any additional costs.
4.3 Adversarial System with Multiple LLMs
In this section, we devise an adversarial system involving multiple LLMs. Given that our method introduces a defense agent against attackers, a natural idea is to then create an attack agent to counter the defender. The attack agent is tasked with generating adversarial examples from purified examples to deceive the target LLM once more. To accomplish this, we design prompts for generating an attack agent $g_{atk}$ by, {mdframed} # Attack Agent Instruction
@@ 4.3 Adversarial System with Multiple LLMs
For example, the original sentence [Purified Example] is classified as [Correct Label]. You should generate a new sentence which is classified as [Incorrect Label]. Input: [Input]. Now, let’s start the attack process and only output the generated sentence.
The prompt structure of attack agent and defense agent is basically the same, although there are some differences in details. The input description of attack agent includes correct label $y$, and the input format is $(x_{in},y)\in D$. The attack instruction is “1. The new sentence should be classified as the opposite of the ‘correct label’. 2. Change at most two letters in the sentence.” Finally, we provide a specific example to help the attack agent better understand the attack task.
Then, we combine the defense agent and attack agent to build an adversarial system, as illustrated in Figure 2. In the adversarial system, the purified examples can be attacked again by the attack agent, and likewise, the adversarial examples can also be purified by the defense agent. They continuously counter each other, much like adversarial training [5].
5 Experiments
We conduct extensive experiments on GLUE datasets to evaluate the effectiveness of the proposed method. Specifically, our method reduces the attack success rate by up to 37.86% with GPT-3.5 and 45.59% with LLAMA-2, respectively.
5.1 Experimental Setup
5.1.1 Datasets
The experiments are conducted on six tasks in GLUE datasets [33], including SST-2 [28], RTE [2], QQP [38], QNLI [23], MNLI-mm, MNLI-m [39]. The detailed descriptions are provided in Appendix A.
5.1.2 Adversarial Attacks
We evaluate our method against PromptAttack [41], which is a powerful attack that combines nine different types of attacks, as illustrated in Table 11. Furthermore, [41] introduces the few-shot (FS) [16] and ensemble (EN) strategies [1] to boost the attack power, details in Appendix B.
5.1.3 Evaluation Metrics
We evaluate the performance of defense methods using two metrics: attack success rate (ASR) and robust accuracy (RA). A lower ASR or higher RA indicates greater model robustness.
5.1.4 Experiment Details
The experiments in this paper are conducted using ‘GPT-3.5-Turbo-0613’ and ‘LLAMA-2-7b’. For GPT-3.5, we purchase OpenAI’s API service and conduct testing experiments with the ‘openai’ package in Python. For LLAMA-2, we deploy it locally on NVIDIA RTX A6000 and utilize the available checkpoint published by MetaAI from HuggingFace.
Table 2: The attack success rate (ASR) defense against PromptAttack-EN and PromptAttack-FS-EN on the GLUE dataset with GPT-3.5.